In a triangle ABC,prove that sin3Asin(B-C)+sin3Bsin(C-A)+sin3Csin(A-B)=0

Grade:Select Grade

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Please approach using

( Sin A/ a) = ( sin B / b ) = ( sin C / c) = k

Expand sin 3A = 3ak - 4 ( ak)^3

If doubt still persists kindly revert for complete solution

Regards